In a family with 4 children, excluding multiple births, what is the probability of having 2 girls and 2 boys, in that order? Assume that a boy is as likely as a girl at each birth. show your work.
Answer by ewatrrr(24785)

P = (1/2)(1/2)(1/2)(1/2) = (1/2)^4 = 1/16